# An Overview of Prototype Formulations for Interpretable Deep Learning

### v3

@@ 4.1 Prototype Losses
The $L_{\mathrm{Sep}}$ function punishes high similarity values between a latent vector ${\bm{z}}_{c}$ of class c and prototypes ${\bm{p}}_{\neg c}$ not belonging to $c$, thereby separating the clusters in latent space. Please note, the original ProtoPNet paper uses a slightly different notation by working with the $L_{2}$ distance, instead of a similarity measure.
@@ 4.2 Multi-Objective Loss Function
4.3 Network Architecture
For our experiments, we use the network architecture proposed by ProtoPNet (Chen et al., 2019), which we deem the basic deep prototype learning architecture for image classification. Figure 4 illustrates this network architecture with the HyperPG prototype module. In the encoding stage, a pretrained image encoder such as a ResNet model (He et al., 2016) serves as model backbone. Then, a convolution neck projects the high dimensional feature map of the backbone into the lower dimensional prototype space. This neck consists of two $1\times 1$ convolution layers with ReLU activation in between followed by a Sigmoid activation.
In the prototype stage, the prototype module computes the similarity scores. The original ProtoPNet uses Euclidian prototypes, which compute the $L_{2}$ distance between the encoded feature maps and the prototype parameters. Figure 4 shows the make up of the HyperPG module: First, the prototype module computes the cosine similarity of the learnable HyperPG anchors $\bm{{\alpha}}$ to the latent vectors produced by the neck. Second, a Density Estimation layer with learnable parameters mean $\mu$ and std $\sigma$ computes the Gaussian PDF over the activations of the previous layer. Future extensions could adapt both components independently, for example by implementing a hyperbolic similarity measure or a multi-modal probability distribution.
The final classification stage produces the output logits based on the prototype module’s similarity scores. ProtoPNet uses a single linear layer due to its inherent interpretability.
4.3.1 Comparison to Recent Works
Our experiments use only the basic ProtoPNet architecture and focus on the effect of different prototype formulations. This section provides a short overview of the required changes to this basic architecture in order to implement the models described in Table 1. LucidPPN (Pach et al., 2024) replaces the encoding stage with a hybrid encoder specializing in color-only and texture-only feature extraction. Deformable ProtoPNet (Donnelly et al., 2021) makes changes towards the input of the prototype module to allow for dynamic spatial arrangements of the prototypes, i.e., “deformations”. ProtoPool (Rymarczyk et al., 2022) makes changes towards the output of the prototype module, by also learning how to dynamically pool the different prototype activations. ProtoTree (Nauta et al., 2021) and ProtoKNN (Ukai et al., 2023) propose to use a different inherently interpretable model for the classification stage. ProtoPFormer (Xue et al., 2024) builds exclusively on a Vision Transformer backbone and uses different network branches for the global information encoded in the class token and the local information in the image tokens.
@@ 5 Experiments
The experiments were performed on CUB-200-2011 (CUB) with 200 bird species (Wah et al., 2011) with ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) backbones. Later experiments also explore the Stanford Cars (Krause et al., 2013) and Oxford Flowers (Nilsback & Zisserman, 2008) datasets. Implementation details such as data preprocessing or model implementation are provided in Appendix D.
Table 2: Mean Top-1 Test Accuracies and Std for different prototype formulations with ResNet50 Backbones averaged over 3 random seeds.
ResNet50 Dense121 ViT-B16
Baseline 79.2 $\pm$ 0.2 75.5 $\pm$ 0.2 78.5 $\pm$ 0.8
Euclidian (ProtoPNet) 35.3 $\pm$ 1.4 33.6 $\pm$ 3.0 9.4 $\pm$ 0.4
Gaussian 61.1 $\pm$ 0.2 45.9 $\pm$ 5.1 4.2 $\pm$ 1.3
Cosine 71.7 $\pm$ 0.4 68.7 $\pm$ 0.5 69.2 $\pm$ 3.9
HyperPG (Ours) 74.3 $\pm$ 0.6 70.7 $\pm$ 0.3 72.9 $\pm$ 0.1
5.1 Quantitative Results
Table 2 reports the mean top-1 accuracy and standard deviation for the tested models on the CUB dataset with backbone models ResNet50 (He et al., 2016), DenseNet121 (Huang et al., 2017) and ViT-B16 (Dosovitskiy, 2020) pretrained on ImageNet. Please note that in contrast to prior work, we do not report scores from model ensembles or using pretrained iNaturalist weights. The black-box baselines perform the best across all experiments, which we see as a theoretical upper bound for the performance. The classical Euclidean prototypes perform the worst across all experiments. A first model improvement can be achieved by switching to probabilistic Gaussian prototypes. The hyperspherical Cosine prototypes further improve the performance. HyperPG combines the advantages of Gaussian and Cosine prototypes, further closing the gap to the upper bound presented by the black-box baselines. Notably, both Euclidian and Gaussian prototypes are unable to learn on the latent features of ViT-B16. We theorize that the attention mechanism of the Transformer architecture produces a near-hyperspherical latent space. This highlights the importance of choosing the right prototype formulation for the used backbone. For example, language aligned models such as (Radford et al., 2021) use the cosine similarity for their alignment. We expect that only hyperspherical prototypes can produce meaningful results in this setting.
@@ 5.1 Quantitative Results
Figure 5: CUB-200-2011 Test Accuracy per Epoch with ResNet50 backbone. Mean and std over 3 random seeds.
Figure 5 shows the test accuracy over epochs for CUB with ResNet50 backbone. The standard Euclidean prototypes take the longest time (200 Epochs) until convergence while also achieving the lowest accuracy. By changing only the prototype formulation to a Gaussian one, the model achieves higher performance and converges at 100 epochs. Both the cosine and HyperPG formulation outperform the other formulations in model accuracy and training time, requiring only 30 training epochs.
5.2 Ablation: Number of Prototypes
We conduct an ablation study over the number of prototypes per class and present the test accuracy scores in Figure 6. Notably, the predictive performance of Euclidean and Gaussian prototypes degrades when more prototypes are added. The prototype formulations with a hyperspherical structure, Cosine and HyperPG prototypes converge to similar performance levels regardless of the number of prototypes per class.
5.3 Ablation: Number of Dimensions
We conduct an ablation study over the number of prototype dimensions and present the test accuracy scores in Figure 7. Hyperspherical prototype formulations such as Cosine and HyperPG are robust regarding the choice of prototype dimensions, although HyperPG performs slightly better with increased dimensionality. The Gaussian prototypes, like HyperPG prototypes, can adapt their variance $\sigma^{2}$ to deal with the increase in dimensionality. The performance of Euclidean prototypes suffers most with changes in dimensionality.
5.4 Ablation: Additional Datasets
We perform additional experiments on two different datasets. Stanford Cars (Cars) with 196 car models (Krause et al., 2013) and Oxford Flowers (Flowers) with 102 different species (Nilsback & Zisserman, 2008). The relative differences in performance seem to hold across the different datasets, with Euclidean prototypes performing the worst. Interestingly, Gaussian prototypes are able to match the performance of the hyperspherical formulations more closely.
Table 3: Top-1 Test Accuracy with Resnet50 Backbone.
Cars Flowers
Baseline 85.2 90.0
Euclidean (ProtoPNet) 36.2 15.0
Gaussian 73.0 74.8
Cosine 79.6 85.5
HyperPG (Ours) 79.0 87.8
@@ 6 Interpretability Analysis
We perform a qualitative analysis of the different prototype formulations with ResNet50 backbone (see Figure 8). One common visualization technique originally proposed by ProtoPNet is a a direct overlay of the prototype activation map (PAM). However, this visualization technique has several shortcomings, which we discuss in more detail in Appendix A. In order to offer an alternative to the PAM visualization, we also present GradCAM (Selvaraju et al., 2017) overlays.
Both Euclidian and HyperPG prototypes do not actually learn class-specific prototypes, in contrast to Gaussian or Cosine prototypes. Interestingly, all prototype formulations seem to activate in similar regions in the test image, but differ with the known training examples. HyperPG could already leverage class-shared prototypes, which would explain its robustness to changes in the number of prototypes per class.
There seems to be no qualitative difference between the visualizations produced with the various prototype formulations using the CNN based backbones. As expected, the ViT backbone (see Figure 9) suffers from the “distraction problem” as observed by ProtoPFormer (Xue et al., 2024). We present additional examples in Appendix A.
Refer to caption Refer to caption
@@
Figure 8: Left: Test Image. Right: 3 closest Training Patches to the Prototype. Top Row: PAM visualization. Bottom Row: GradCAM visualization. R50: Resnet50 Backbone. ViT: ViT-B16 Backbone.
@@ 7 Conclusion
This work provides comprehensive overview of different prototype formulations, with both probabilistic and point-based measurements, in an Euclidean or hyperspherical latent space. We introduce HyperPG, a new prototype representation based on Gaussian distributions on the surface of a hypersphere. Our experiments showcase the differences in performance and robustness to hyperparameters across different datasets and backbone architectures for these prototype formulations. Our results demonstrate the performance advantage of hypersherical prototypes and probabilistic prototypes in contrast to the commonly used Euclidean formulation, with no loss to the inherent interpretability.
Future work could explore further refinements to the probabilistic prototype formulations, including adaptive mechanisms for prototype selection or the integration of additional probabilistic models such as Mixture Models or Bayesian approaches like MGProto (Wang et al., 2024b) to further enhance interpretability and performance.
@@ A.1 Potential Problems with Prototype Activation Maps
Appendix B Adapting HyperPG to other Probability Distributions
@@ B.1 Cauchy Distribution
B.2 Truncated Distributions
The cosine similarity is defined only on the interval $[-1,1]$. This makes it attractive to also use truncated probability distributions, which are also only defined on this interval. The truncation imposes a limit on the range of the PDF, thereby limiting the influence of large values for the distribution’s $\sigma$ or $\gamma$ parameter, respectively. The truncated Gaussian pdf $\mathcal{T}_{\mathrm{Gauss}}$ requires the cumulative probability function $\mathcal{G}$ and error function $f_{\mathrm{err}}$, and is defined as
[…]
@@ B.2 Truncated Distributions
with lower bound $a$ and upper bound $b$, e.g., for the cosine similarity $a=-1$ and $b=1$. Similarly, the truncated Cauchy distribution can be applied, which is defined as
@@
B.3 von Mises-Fisher Distribution
@@
with mean vector $\bm{{\alpha}}$, scalar concentration parameter $\kappa$ and normalization constant $C_{d}(\kappa)$. The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^1] Tensorflow API Documentation - Accessed 2025-01-10 only supports the vMF distribution for $D\leq 5$.
@@
B.4 Fisher-Bingham Distribution

### v4

@@ 4.2 Prototype Losses
The $L_{\mathrm{Sep}}$ function punishes high similarity values between a latent vector ${\bm{z}}_{c}$ of class c and prototypes ${\bm{p}}_{\neg c}$ not belonging to $c$, thereby separating the clusters in latent space.Please note, the original ProtoPNet implementation is optimized for minimizing the $L_{2}$ distance, instead of maximizing the similarity measure. Using the $L_{2}$ similarity would require tuning of the associated hyperparameters.
@@ 4.3 Multi-Objective Loss Function
Table 2: Mean Top-1 Test Accuracies (%) CUB-200-2011.
ProtoPNet Optim. Simplified Optim. Add. Datasets (R50)
Method R50 D121 R50* D121* ViT* Cars* Flowers*
Baseline $79.2$ p m 0.2 $75.5$ p m 0.2 $79.2$ p m 0.2 $75.5$ p m 0.2 $78.5$ p m 0.8 $85.2$ $90.0$
Euclidean (PPN) $73.9$ p m 0.4 $76.4$ p m 1.2 $61.4$ p m 1.4 $57.8$ p m 3.0 $23.4$ p m 0.4 $75.4$ $70.7$
Gaussian $6.4$ p m 0.3 $75.7$ p m 0.2 $61.1$ p m 0.2 $45.9$ p m 5.1 $4.2$ p m 1.3 $73.0$ $74.8$
Cosine $78.5$ p m 0.2 $80.4$ p m 0.2 $71.7$ p m 0.4 $68.7$ p m 0.5 $69.2$ p m 3.9 $79.6$ $85.5$
HyperPG (Ours) $77.8$ p m 0.2 $78.7$ p m 0.1 $74.3$ p m 0.6 $70.7$ p m 0.3 $72.9$ p m 0.1 $79.0$ $87.8$
*Without warm-up epochs and training optimizations. R50: ResNet50, D121: DenseNet121
4.4 Network Optimization
ProtoPNet proposes a sophisticated network optimization scheme to achieve its high performance. Initially, for five warm-up epochs, only the convolution neck and prototype module are trained using Adam (Kingma and Ba, 2015) with learning rate $3\text{\times}{10}^{-3}$, while the classification layer uses $1\text{\times}{10}^{-4}$. During the warm-up phase, the pretrained image encoder remains frozen. After the warm-up phase, a new Adam optimizer trains the encoder with learning rate $1\text{\times}{10}^{-4}$, while the neck and prototype module continue at $3\text{\times}{10}^{-3}$. A scheduler reduces learning rates by $1\text{\times}{10}^{-1}$ every 5 epochs.
A preliminary Bayesian hyperparameter optimization using Weights and Biases (Biewald, 2020) yielded no significant improvements, suggesting the original ProtoPNet implementation is already highly optimized.
@@ 5 Experiments
Table 2 reports the mean top-1 accuracy and standard deviation for the tested models across three random seeds. Unlike prior work, we use only ImageNet weights, omitting model ensembles and pretrained iNaturalist weights. We first conduct image classification experiments on CUB-200-2011 (Wah et al., 2011) with the full architecture and training implementation as proposed by ProtoPNet (Chen et al., 2019), then with a simplified network optimization, and finally on two additional image classification datasets. Implementation details such as data preprocessing or model implementation are provided in Appendix B.
5.1 ProtoPNet Optimization
We experiment on the CUB dataset using ResNet50 (He et al., 2016) and DenseNet121 (Huang et al., 2017) backbones pretrained on ImageNet with ProtoPNet’s full optimization scheme (Chen et al., 2019). We also test a blackbox baseline without a prototype module. Table 2 (ProtoPNet Optim.) presents the top-1 test accuracies on CUB-200-2011.
In this experiment, we evaluate different prototype formulations without adjusting hyperparameters. Cosine prototypes perform best, even outperforming the Dense121 blackbox baseline. While probabilistic formulations perform slightly worse than their point-based counterparts, Gaussian prototypes show high sensitivity to hyperparameters and backbone selection. They fail to learn meaningful representations with the ResNet50 backbone.
5.2 Simplified Optimization
To evaluate the robustness and practical applicability of different prototype formulations, we test them under a drastically simplified training scheme that removes ProtoPNet’s multi-stage optimization, warm-up epochs, and learning rate scheduling in favor of a single AdamW optimizer (learning rate $1\text{\times}{10}^{-4}$, no scheduling) for the entire model. Table 2 (Simplified Optim.) shows the top-1 accuracies on CUB-200-2011. The Euclidean prototype formulation shows heavily reduced performance, suggesting that the original ProtoPNet implementation was highly finetuned. While hyperspherical prototypes also perform worse, remain competitive with the fully-optimized Euclidean ProtoPNet.
@@ 5.2 Simplified Optimization
Figure 5: CUB-200-2011 Test Accuracy per Epoch with ResNet50 backbone and simplified optimization scheme. Mean and std over 3 random seeds.
Figure 5 illustrates the Top-1 Test Accuracy on CUB-200-2011 over training epochs. With the simplified training scheme, the hyperspherical prototype formulations converge much faster than $L_{2}$-based formulations. Cosine and HyperPg converge within 20 training epochs, whereas the Euclidean prototypes require over 200 epochs.
5.3 Ablation: Transformer Backbone
Using the simplified optimization scheme, we test a Vision Transformer backbone. Notably, both Euclidean and Gaussian prototypes fail to learn a latent structure from ViT-B16 features. We hypothesize that the Transformer’s attention mechanism produces a near-hyperspherical latent space. This underscores the importance of matching prototype formulation to backbone architecture. For instance, language aligned models like CLIP (Radford et al., 2021) use the cosine similarity for modality alignment. We anticipate that only hyperspherical prototypes can produce meaningful results when using CLIP as a feature extractor.
5.4 Ablation: Datasets
We apply our simplified training scheme to two additional datasets: Stanford Cars (Cars) with 196 car models (Krause et al., 2013) and Oxford Flowers (Flowers) with 102 species (Nilsback and Zisserman, 2008). Table 2 (Add. Datasets) shows top-1 test accuracies. The performance pattern among prototype formulations remains consistent across datasets, with Euclidean prototypes consistently underperforming.
@@ 6 Interpretability Analysis
We qualitatively analyze different prototype formulations with a ResNet50 backbone and the simplified optimization scheme. ProtoPNet originally proposed directly overlaying the prototype activation map (PAM). However, this visualization technique has several shortcomings, discussed in Appendix A. As an alternative, we also present GradCAM (Selvaraju et al., 2017) visualizations in Appendix A.
Figure 6 shows PAMs for different prototype formulations on the same test image (left), with highest activations marked by yellow rectangles. Following the prototypical concepts approach (Ma et al., 2023), we present the two most highly activated training patches, as not all prototype formulations support a “push” onto training samples.
Interestingly, all prototype formulations activate in similar test image regions but differ in their associated training examples. The Euclidean prototype formulation seems to produce less visually consistent prototype associations than the other formulations using the simplified optimization scheme.
We observe no qualitative differences between visualizations produced by various prototype formulations using CNN-based backbones (see Appendix A). While ViT backbone performs well with cosine prototypes, it exhibits the ”distraction problem” (Figure 7) first observed by ProtoPFormer (Xue et al., 2024). For an extensive overview of prototype examples, GradCAM activations, and additional backbones, refer to Appendix A.
@@
Figure 6: Left: Test Image. Right: Three closest Training Patches to the Prototype.
@@ 7 Conclusion
This work provides a comprehensive overview of different prototype formulations, with both probabilistic and point-based measurements, in Euclidean or hyperspherical latent spaces. We introduce HyperPG, a new prototype representation based on Gaussian distributions on the surface of a hypersphere.
Our experiments demonstrate substantial differences in robustness across prototype formulations. Hyperspherical prototypes maintain competitive performance under drastically simplified training (single optimizer, no warm-up epochs or learning rate scheduling), whereas Euclidean prototypes show severe performance degradation under identical conditions. This indicates that the original ProtoPNet’s strong performance depends heavily on extensive hyperparameter tuning, whereas hyperspherical formulations offer greater optimization stability. This has practical implications for deployment: hyperspherical prototypes provide improved performance with reduced sensitivity to training configuration, making prototype-based models more accessible for practitioners without extensive hyperparameter search resources.
Our results show the advantages of hyperspherical and probabilistic prototypes over the commonly used Euclidean formulation, with no loss to inherent interpretability. Future work could explore further refinements to probabilistic prototype formulations, including adaptive mechanisms for prototype selection or the integration of additional probabilistic models such as Mixture Models or Bayesian approaches like MGProto (Wang et al., 2024b).
Acknowledgements
The research presented in this paper was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 448648559.
@@ Appendix A Extended Interpretability Analysis
Figure 8: DenseNet121 Euclidean. Top: PAM. Bottom: GradCAM
Figure 9: DenseNet121 Gaussian. Top: PAM. Bottom: GradCAM
Figure 10: DenseNet121 Cosine. Top: PAM. Bottom: GradCAM
Figure 11: DenseNet121 HyperPG. Top: PAM. Bottom: GradCAM
Figure 12: ViT-B16 Euclidean. Top: PAM. Bottom: GradCAM
Figure 13: ViT-B16 Gaussian. Top: PAM. Bottom: GradCAM
Figure 14: ViT-B16 Cosine. Top: PAM. Bottom: GradCAM
Figure 15: ViT-B16 HyperPG. Top: PAM. Bottom: GradCAM
@@ A.1 Potential Problems with Prototype Activation Maps
Appendix B Implementation Details
Table 3: Hyperparameters used in our experiments
Hyperparameter PPN HyperPG
Data and Model Configuration
Image size 224 224
Number of prototypes per class 10 10
Training batch size 80 48
Testing batch size 100 48
Training Schedule
Number of training epochs 20 (offline augmentation $\times 30$) 200 (online augmentation)
Number of warmup epochs 5 N.A.
Warm Optimizer Learning Rates
Add-on layers $3\times 10^{-3}$ N.A.
Prototype vectors $3\times 10^{-3}$ N.A.
Optimizer Learning Rates
Features $\times 10^{-4} 1×1$$\times 10^{-4}$
Add-on layers $3\times 10^{-3}$ $1\times 10^{-4}$
Prototype vectors $3\times 10^{-3}$ $1\times 10^{-4}$
Joint learning rate step size 5 N.A.
Final Layer Training
Last layer optimizer learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$
Compute Resources
GPU Model NVidia RTX 4090 NVidia RTX 3070Ti
GPU VRAM 24 GB 8 GB
Our experiments can be divided into two implementations: “PPN”, built on the original ProtoPNet (Chen et al., 2019) implementation published on Github [^2] https://github.com/cfchen-duke/ProtoPNet with a sophisticated network optimization scheme. And “HyperPG”, our own implementation [^3] https://github.com/LiXiling/prob-proto built from scratch, originally designed for proposing and testing the HyperPG prototypes.
B.1 Data Preprocessing
B.1.1 PPN: Data Preprocessing
ProtoPNet uses an offline data augmentation process, resulting in 30 training images per class and epoch. The input images are cropped to the bounding box annotations. The images are augmented by applying Rotation XOR Skewing XOR Shearing, followed by a 50% chance of left-right flipping.
B.1.2 HyperPG: Data Preprocessing
B.2 Hyperparameters
For both implementations we provide the full hyperparameters in Table 3. PPN uses the exact hyperparameters as used by ProtoPNet according to the publically available Github implementation.
Our HyperPG implementation does not use a training scheme with warm-up epochs or learn-rate schedulings, making several hyperparameters not applicable, marked as N.A. The batchsizes are optimized for the different hardware configurations.
B.3 Compute Resources
The PPN experiments were performed on a computing node with four NVidia RTX 4090 GPUs (24 GB VRAM each), with each model accessing one GPU. Please note, that the batch size could have been further increased on this hardware configuration. On CUB-200-2011 the training duration was roughly 20 minutes per epoch, with no meaningful difference between prototype formulations.
The HyperPG experiments with simplified training and online data augmentation were performed on a desktop workstation with a single NVIDIA RTX 3070Ti GPU (8 GB VRAM). On CUB-200-2011 the training duration was roughly 40 seconds per epoch, with no meaningful difference between prototype formulations.
We estimate 160 GPU hours for running the PPN experiments plus 120 GPU hours for the HyperPG experiments. With an additional, generous 15% overhead estimation for debugging, analyses, etc. we feel confident that this work took less than 320 GPU hours to develop.
B.4 Model Implementation
Appendix C Additional Ablations
C.1 Ablation: Number of Prototypes
We conduct an ablation study over the number of prototypes per class using the simplified training scheme and present the test accuracy scores in Figure 16. Notably, the predictive performance of Euclidean and Gaussian prototypes degrades when more prototypes are added. The prototype formulations with a hyperspherical structure, Cosine and HyperPG prototypes converge to similar performance levels regardless of the number of prototypes per class.
C.2 Ablation: Number of Dimensions
We conduct an ablation study over the number of prototype dimensions using the simplified training scheme and present the test accuracy scores in Figure 17. Hyperspherical prototype formulations such as Cosine and HyperPG are robust regarding the choice of prototype dimensions, although HyperPG performs slightly better with increased dimensionality. The Gaussian prototypes, like HyperPG prototypes, can adapt their variance $\sigma^{2}$ to deal with the increase in dimensionality. The performance of Euclidean prototypes suffers most with changes in dimensionality.
Appendix D Adapting HyperPG to other Probability Distributions
@@ D.1 Cauchy Distribution
D.2 Truncated Distributions
The cosine similarity is defined only on the interval [1,1][-1,1]. This makes it attractive to also use truncated probability distributions, which are also only defined on this interval. The truncation imposes a limit on the range of the PDF, thereby limiting the influence of large values for the distribution’s $\sigma$ or $\gamma$ parameter, respectively. The truncated Gaussian pdf $\mathcal{T}_{\mathrm{Gauss}}$ requires the cumulative probability function $\mathcal{G}$ and error function $f_{\mathrm{err}}$, and is defined as
$f_{\mathrm{err}}(x)$ $=\frac{2}{\sqrt{\pi}}\int_{0}^{x}\exp\left(-z^{2}\right)dz,$ (10)
@@ D.2 Truncated Distributions
with lower bound $a$ and upper bound $b$, e.g., for the cosine similarity a=1a=-1 and $b=1$. Similarly, the truncated Cauchy distribution can be applied, which is defined as
@@
D.3 von Mises-Fisher Distribution
@@
with mean vector $\bm{{\alpha}}$, scalar concentration parameter $\kappa$ and normalization constant $C_{d}(\kappa)$. The normalization constant $C_{d}(\kappa)$ is a complex function and difficult to compute for higher dimensions, which is why, for example, Tensorflow [^4] Tensorflow API Documentation - Accessed 2026-01-07 only supports the vMF distribution for $D\leq 5$.
@@
D.4 Fisher-Bingham Distribution